# bibdata.bib

### v1

@@ V-A Experimental Setup
Benchmarks used in our experiments are listed in Table I, which includes circuit names, PI/PO numbers, AIG size, AIG depth, circuit area, and circuit delay. They are the benchmarks used in MECALS [meng2023mecals], EPFL arithmetic benchmarks [lsi2024epfl], and ISCAS85 benchmarks [hansen1999unveiling]. The initial AIGs have been well optimized to ensure as little redundancy as possible. These AIGs are then used as input to our ALS flow and those ALS flows for comparison. The baseline methods are the MECALS [meng2023mecals] and MUSCAT [witschen2022muscat] methods. MECALS is a state-of-the-art ALS method, in which the maximum error checking problem is converted into a SAT sweeping problem. MUSCAT is another state-of-the-art method that converts the ALS problem under maximum error constraint into a MUS problem and solves it using a MUS solver. MECALS can handle both constant [shin2011new] and SASIMI [venkataramani2013substitute] LACs used in the experiments, while MUSCAT can only handle the constant LAC.
TABLE II: Comparison of our method with the state-of-the-art methods under the MaxED constraint. Bold entries indicate that our method outperforms MECALS and MUSCAT. N/A means MUSCAT cannot obtain the final approximate circuit in 24 hours.
Circuit MaxED bound Area ratio Delay ratio Runtime/s
Ours MECALS MUSCAT Ours MECALS MUSCAT Ours MECALS MUSCAT
absdiff 1 64.0% 65.2% 86.2% 107.2% 108.9% 94.6% 0.2 0.4 1.1
3 59.7% 69.2% 84.7% 95.1% 106.2% 105.7% 1.7 0.6 1.1
add8 1 76.0% 92.5% 95.0% 77.3% 88.1% 101.6% 0.1 0.1 0.7
3 72.2% 81.7% 85.5% 76.7% 95.7% 87.0% 0.7 0.2 1.0
add32 9 70.3% 71.9% 84.1% 95.0% 99.6% 76.3% 28 50 23
97 63.0% 62.5% 82.6% 84.9% 86.0% 83.7% 114 84 28
binsqrd 3 78.4% 79.9% 99.9% 95.2% 93.8% 100.0% 15 1432 377
12 76.9% 78.6% 97.2% 95.2% 93.9% 99.7% 271 8947 7230
buttfly 10 79.4% 86.9% 97.7% 81.6% 91.8% 99.3% 15 3.6 2.2
111 74.6% 85.3% 94.9% 101.1% 83.7% 99.3% 56 5.8 2.3
mac 1 87.1% 86.6% 96.6% 94.4% 95.8% 99.6% 0.3 1.0 1.9
2 85.1% 82.8% 89.4% 94.4% 91.5% 87.4% 0.8 4.5 4.5
mult8 3 74.1% 75.1% 98.2% 81.1% 82.0% 100.0% 1.3 78 949
9 72.5% 73.1% 96.3% 85.9% 81.9% 100.0% 7.7 111 7225
Average of above 73.8% 77.9% 92.0% 90.4% 92.8% 95.3% 37 766 1132
mult16 9 98.3% 99.7% N/A 98.1% 100.7% N/A 9.3 154 N/A
84 93.7% 95.7% N/A 92.1% 97.1% N/A 83 368 N/A
mult32 84 98.0% 99.6% N/A 97.5% 102.0% N/A 203 997 N/A
7131 93.9% 98.6% N/A 100.9% 93.8% N/A 1544 3364 N/A
Average of all 78.7% 82.5% N/A 91.9% 94.0% N/A 131 867 N/A
@@ V-B1 Comparison with State-of-the-Art Methods
We compare our ALS flow with MECALS and MUSCAT. The tested benchmarks are those used in MECALS. We run the open-source codes of MECALS and MUSCAT on the same platform for fair comparison. The MaxED bounds are set to $\lfloor 2^{0.1O}\rfloor$ and $\lfloor 2^{0.2O}\rfloor$ for each benchmark, where $O$ is the number of POs of the benchmark, and the function $\lfloor x\rfloor$ gives the greatest integer less than or equal to $x$. The constant and SASIMI LACs are used in our method and MECALS, while MUSCAT uses its only supported constant LAC.
@@
Table II compares the area ratio, delay ratio, and runtime of our flow, MECALS, and MUSCAT under the MaxED constraint. The bold entries indicate that our method outperforms both MECALS and MUSCAT, and we use the same highlighting in the following tables. We can see that our method achieves the smallest area ratios for most benchmarks. For the first 7 smaller benchmarks, our method achieves an average area ratio of $73.8\%$ and an average delay ratio of $90.4\%$. Compared to MECALS, our method reduces area and delay by $4.1\%$ and $2.4\%$  respectively, on average. Compared to MUSCAT, our method reduces area and delay by 18.2% and 4.9%, respectively, on average. Moreover, MUSCAT cannot obtain the final approximate circuit in 24 hours for the benchmarks mult16 and mult32, while our method and MECALS can. Over all benchmarks, our method reduces area and delay by $3.8\%$ and $2.1\%$, respectively, than MECALS on average. For the benchmarks absdiff, add8, mult8, and mult16, our method achieves both smaller area and delay ratios than MECALS. However, for the benchmark add32 under the MaxED bound of $97$ and the benchmark mac, our method is worse than MECALS in terms of the area ratio. One possible reason is as follows. Both our flow and MECALS iteratively simplify the circuit. In each iteration, our flow selects multiple promising LACs (see Section IV-C2), while MECALS only selects one. This difference may lead to better performance of MECALS on some benchmarks. However, our flow reduces more area than MUSCAT for all benchmarks. This is because our flow can handle more complex LACs than MUSCAT, which can achieve better approximate circuits.
@@
Our method is more efficient than MECALS and MUSCAT. Over the first 7 benchmarks, our method is 20.7×20.7$\times$20.7 × faster than MECALS and 30.6×30.6$\times$30.6 × faster than MUSCAT on average. Over all benchmarks, our method speeds up by $\times6$ on average compared to MECALS. Note that our method is slower than MECALS for some benchmarks. For the small benchmarks absdiff and add8, the total runtime of our method is within $2$ seconds, so the runtime difference is negligible. For the benchmark buttfly, compared to MECALS, our method takes more time but dramatically reduces the area, which is worth the trade-off. For the benchmark add32 under the MaxED bound of $97$, our method consumes more time than MECALS and MUSCAT, while the area and delay ratios are still competitive.
@@
TABLE III: Comparison of our ALS flow with the MECALS method on the EPFL arithmetic benchmarks under the MaxED constraint. Bold entries indicate that our method outperforms MECALS. N/A means that MECALS cannot obtain the final approximate circuit in 24 hours.
@@
Circuit MaxED bound Area ratio Delay ratio Runtime/s
Ours MECALS Ours MECALS Ours MECALS
add128 7.6×$\times 10^{3}$ 86.9% 94.2% 105.4% 96.4% 7.2 474
5.8×$\times 10^{7}$ 77.6% 84.1% 99.9% 100.1% 11 3264
bar 7.1×$\times 10^{3}$ 97.5% 97.9% 99.9% 100.6% 2.2 412
5.1×$\times 10^{7}$ 96.2% 95.7% 100.2% 100.4% 2.6 753
max 8.2×$\times 10^{3}$ 94.2% 94.2% 82.4% 82.4% 273 6362
6.7×$\times 10^{7}$ 93.1% 93.2% 80.8% 80.9% 303 8580
mult64 7.1×$\times 10^{3}$ 96.1% 99.1% 100.6% 98.5% 4347 45437
5.1×$\times 10^{7}$ 95.7% 98.2% 102.4% 99.2% 1996 78970
square 7.1×$\times 10^{3}$ 92.9% 99.4% 93.3% 98.9% 1504 53925
Average of above 92.2% 95.1% 96.1% 95.3% 938 22020
square 5.1×$\times 10^{7}$ 93.0% N/A 89.8% N/A 13903 N/A
log2 9.0×$\times 10^{0}$ 94.1% N/A 104.8% N/A 37143 N/A
8.4×$\times 10^{1}$ 93.9% N/A 103.7% N/A 36754 N/A
sin 5.0×$\times 10^{0}$ 94.9% N/A 110.2% N/A 1195 N/A
3.2×$\times 10^{1}$ 79.7% N/A 96.7% N/A 22293 N/A
sqrt 8.4×$\times 10^{1}$ 81.2% N/A 81.3% N/A 400 N/A
7.1×$\times 10^{3}$ 62.6% N/A 62.9% N/A 363 N/A
Average of all 89.3% N/A 94.6% N/A 7531 N/A
@@ V-B2 Experiments on EPFL Benchmarks
To show the scalability of our ALS flow, we test it on the large EPFL arithmetic benchmarks. The MUSCAT method cannot handle them, so we do not compare our method with MUSCAT and we only compare our method with MECALS. Similarly, the MaxED bounds are set to $\lfloor 2^{0.1O}\rfloor$ and $\lfloor 2^{0.2O}\rfloor$ for each benchmark, where $O$ is the PO number of the benchmark. To accelerate our flow, we first apply the constant LACs to quickly simplify the circuit, followed by the SASIMI LACs for further simplification. Specifically, the constant LACs are first applied to the circuit POs (like truncation) until the MaxED bound is reached and then to the internal nodes of the circuit until the MaxED bound is reached. Finally, the SASIMI LACs are applied until the MaxED bound is reached.
@@
Table III compares the area ratio, delay ratio, and runtime of our flow and MECALS on the EPFL arithmetic benchmarks under the MaxED constraint. We can see that our method can handle all benchmarks in the table with an average runtime of $7531$ seconds, while MECALS cannot handle the benchmarks log2, sin, sqrt, and square (under $5.1\times 10^{7}$ MaxED bound) in 24 hours. For the benchmarks that both our method and MECALS can handle (the top part of the table), our method further reduces the area by $2.9\%$ on average with a small delay overhead, while it is 23.5×23.5$\times$23.5 × faster than MECALS on average. Notably, for the benchmark add128, under the two MaxED bounds of $7.6\times 10^{3}$ and $5.8\times 10^{7}$, our method speeds up by 65.8×65.8$\times$65.8 × and 296.7×296.7$\times$296.7 ×, respectively, compared to MECALS and reduces area by $7.3\%$ and $6.5\%$, respectively. Only for the benchmark bar under the MaxED bound of $5.1\times 10^{7}$, our method is slightly worse than MECALS in terms of the area ratio. However, our method is far more efficient than MECALS with a competitive delay ratio.
@@
TABLE IV: Comparison of our ALS flow with the MECALS method on the ISCAS benchmarks under the MaxHD constraint. Bold entries indicate that our method outperforms MECALS.
@@
Circuit MaxHD bound Area ratio Delay ratio Runtime/s
Ours MECALS Ours MECALS Ours MECALS
c880 2 88.3% 97.3% 100.5% 105.6% 0.9 1.9
5 67.5% 92.4% 113.1% 106.8% 1.1 2.8
c1355 3 10.1% 10.4% 3.4% 3.4% 0.6 73
6 9.1% 10.8% 3.4% 3.4% 0.4 17
c1908 2 89.9% 97.3% 89.5% 101.1% 1.0 5.0
5 73.6% 65.5% 84.3% 86.6% 1.5 53
c2670 14 64.2% 84.9% 62.9% 99.6% 29 26
28 42.3% 41.2% 36.9% 54.4% 32 80
c3540 2 93.8% 95.5% 92.8% 109.1% 1.3 64
4 92.1% 95.6% 98.7% 102.8% 1.6 105
c5315 12 92.7% 96.5% 104.2% 103.5% 25 158
24 84.6% 90.5% 93.8% 99.6% 34 877
c7552 10 41.2% 95.8% 77.6% 99.8% 16 193
21 32.7% 84.2% 68.2% 100.1% 29 627
Average 63.0% 75.6% 73.5% 84.0% 12 163

### v2

@@ V-A Experimental Setup
The benchmarks used in our experiments are listed in Table I, which includes circuit names, PI/PO numbers, AIG size, AIG depth, circuit area, circuit delay, and circuit power. They are the benchmarks used in MECALS [17], EPFL arithmetic benchmarks [9], and ISCAS85 benchmarks [11]. The initial AIGs have been well optimized to ensure as little redundancy as possible. These AIGs are then used as input to our ALS flow and those ALS flows for comparison. The baseline methods are the MECALS [17] and MUSCAT [42] methods. MECALS is a state-of-the-art ALS method, in which the maximum error checking problem is converted into a SAT sweeping problem. MUSCAT is another state-of-the-art method that converts the ALS problem under maximum error constraint into a MUS problem and solves it using a MUS solver. MECALS can handle both constant [31] and SASIMI [37] LACs, while MUSCAT only supports constant LACs. In our experiments, for a circuit with $N$ nodes, our method and MECALS consider $2N$ constant LACs and $O(N^{2})$ SASIMI LACs, while MUSCAT only considers the $2N$ constant LACs.
TABLE II: Comparison of our method with the state-of-the-art methods under the MaxED constraint. Bold entries indicate the smallest area, delay, and power ratios, or the shortest runtime. N/A means MUSCAT cannot obtain final approximate circuits in 24 hours.
Circuit MaxED bound Area ratio Delay ratio Power ratio Runtime/s
Ours MECALS MUSCAT Ours MECALS MUSCAT Ours MECALS MUSCAT Ours MECALS MUSCAT
absdiff 1 64.0% 65.2% 86.2% 107.2% 108.9% 94.6% 68.1% 65.2% 83.7% 0.2 0.4 1.1
3 59.7% 69.2% 84.7% 95.1% 106.2% 105.7% 59.5% 62.4% 70.3% 1.7 0.6 1.1
add8 1 76.0% 92.5% 95.0% 77.3% 88.1% 101.6% 81.4% 92.3% 94.3% 0.12 0.11 0.7
3 72.2% 81.7% 85.5% 76.7% 95.7% 87.0% 76.7% 80.7% 85.6% 0.7 0.2 1.0
add32 9 70.3% 71.9% 84.1% 95.0% 99.6% 76.3% 77.6% 79.6% 100.0% 28 50 23
97 63.0% 62.5% 82.6% 84.9% 86.0% 83.7% 69.5% 68.5% 82.6% 114 84 28
binsqrd 3 78.4% 79.9% 99.9% 95.2% 93.8% 100.0% 83.1% 84.0% 100.4% 15 1432 377
12 76.9% 78.6% 97.2% 95.2% 93.9% 99.7% 81.6% 82.7% 98.2% 271 8947 7230
buttfly 10 79.4% 86.9% 97.7% 81.6% 91.8% 99.3% 81.5% 83.6% 93.1% 15 3.6 2.2
111 74.6% 85.3% 94.9% 101.1% 83.7% 99.3% 77.3% 78.8% 86.9% 56 5.8 2.3
mac 1 87.1% 86.6% 96.6% 94.4% 95.8% 99.6% 92.7% 91.3% 92.4% 0.3 1.0 1.9
2 85.1% 82.8% 89.4% 94.4% 91.5% 87.4% 88.3% 87.8% 82.4% 0.8 4.5 4.5
mult8 3 74.1% 75.1% 98.2% 81.1% 82.0% 100.0% 75.1% 72.4% 98.6% 1.3 78 949
9 72.5% 73.1% 96.3% 85.9% 81.9% 100.0% 76.3% 70.8% 97.3% 7.7 111 7225
Average of above 73.8% 77.9% 92.0% 90.4% 92.8% 95.3% 77.8% 78.6% 90.4% 37 766 1132
mult16 9 98.3% 99.7% N/A 98.1% 100.7% N/A 99.0% 99.9% N/A 9.3 154 N/A
84 93.7% 95.7% N/A 92.1% 97.1% N/A 94.5% 96.0% N/A 83 368 N/A
mult32 84 98.0% 99.6% N/A 97.5% 102.0% N/A 98.2% 99.1% N/A 203 997 N/A
7131 93.9% 98.6% N/A 100.9% 93.8% N/A 94.7% 97.8% N/A 1544 3364 N/A
Average of all 78.7% 82.5% N/A 91.9% 94.0% N/A 82.0% 82.9% N/A 131 867 N/A
@@ V-B1 Comparison with State-of-the-Art Methods
We compare our ALS flow with MECALS and MUSCAT. The tested benchmarks are those used in MECALS. We run the open-source codes of MECALS and MUSCAT on the same platform for fair comparison. The MaxED bounds are set to $\lfloor 2^{O/10}\rfloor$ and $\lfloor 2^{O/5}\rfloor$ for each benchmark, where $O$ is the number of POs of the benchmark, and the function $\lfloor x\rfloor$ gives the greatest integer less than or equal to $x$. In each iteration of our flow, the LAC generation step (see Fig. 2) produces both constant and SASIMI LACs.
@@
Table II compares the area ratio, delay ratio, power ratio, and runtime of our flow, MECALS, and MUSCAT under the MaxED constraint. The bold entries indicate the smallest area ratio, delay ratio, power ratio, and runtime among the three methods, and we use the same highlighting in the following tables. We can see that our method achieves the smallest area ratios for most benchmarks. For the first 7 smaller benchmarks, our method achieves an average area, delay, and power ratio of $73.8\%$, $90.4\%$, and $77.8\%$, respectively. Compared to MECALS, our method reduces area, delay, and power by $4.1\%$, $2.4\%$  and 0.8% on average. Compared to MUSCAT, our method reduces area, delay, and power by 18.2%, 4.9%, and 12.6% on average. Moreover, MUSCAT cannot obtain the final approximate circuit in 24 hours for the benchmarks mult16 and mult32, while our method and MECALS can. Over all benchmarks, our method reduces area, delay, and power by $3.8\%$, $2.1\%$, and 0.9% on average compared to MECALS. Note that for the benchmarks add8 and mult16, our method achieves smaller area, delay, and power ratios than MECALS for both MaxED bounds. However, for the benchmark add32 under the MaxED bound of $97$ and the benchmark mac, our method is worse than MECALS in terms of the area ratio. One possible reason is as follows. Both our flow and MECALS iteratively simplify the circuit. In each iteration, our flow selects multiple promising LACs (see Section IV-C2), while MECALS only selects one. This difference may lead to better performance of MECALS on some benchmarks. However, our flow reduces more area than MUSCAT for all benchmarks. This is because our flow can handle more complex LACs than MUSCAT, which can achieve better approximate circuits.
@@
Our method is more efficient than MECALS and MUSCAT. Over the first 7 benchmarks, our method is $20.7\times$ faster than MECALS and $30.6\times$ faster than MUSCAT on average. Over all benchmarks, our method speeds up by $6.6\times$ on average compared to MECALS. Note that our method is slower than MECALS for some benchmarks. For the small benchmarks absdiff and add8, the total runtime of our method is within $2$ seconds, so the runtime difference is negligible. For the benchmark buttfly, compared to MECALS, our method takes more time but dramatically reduces the area, which is worth the trade-off. For the benchmark add32 under the MaxED bound of $97$, our method consumes more time than MECALS and MUSCAT, while the area and delay ratios are still competitive. Additionally, across all benchmarks in Table II, our method finishes after an average of 5.8 iterations. The average number of LACs applied per iteration is 7.4, obtained by dividing the total number of applied LACs across all iterations and benchmarks over the total number of iterations across all benchmarks. In contrast, MECALS needs an average of 8.3 iterations with only one LAC applied per iteration. Applying multiple LACs per iteration reduces the number of iterations, thereby shortening the overall runtime.
@@
TABLE III: Comparison of our ALS flow with the MECALS method on the EPFL arithmetic benchmarks under the MaxED constraint. Bold entries indicate smaller area, delay, and power ratios, or shorter runtime. N/A means that MECALS cannot obtain the final approximate circuit in 24 hours.
@@
Circuit MaxED bound Area ratio Delay ratio Power ratio Runtime/s
Ours MECALS Ours MECALS Ours MECALS Ours MECALS
add128 7.6E+03 86.9% 94.2% 105.4% 96.4% 86.6% 92.0% 7.2 474
5.8E+07 77.6% 84.1% 99.9% 100.1% 77.1% 81.2% 11 3264
bar 7.1E+03 97.5% 97.9% 99.9% 100.6% 95.1% 96.0% 2.2 412
5.1E+07 96.2% 95.7% 100.2% 100.4% 90.8% 90.3% 2.6 753
max 8.2E+03 94.2% 94.2% 82.4% 82.4% 90.5% 90.1% 273 6362
6.7E+07 93.1% 93.2% 80.8% 80.9% 87.5% 87.9% 303 8580
mult64 7.1E+03 96.1% 99.1% 100.6% 98.5% 98.3% 101.0% 4347 45437
5.1E+07 95.7% 98.2% 102.4% 99.2% 97.8% 99.6% 1996 78970
square 7.1E+03 92.9% 99.4% 93.3% 98.9% 96.3% 99.1% 1504 53925
Average of above 92.2% 95.1% 96.1% 95.3% 91.1% 93.0% 938 22020
square 5.1E+07 93.0% N/A 89.8% N/A 95.7% N/A 13903 N/A
log2 9.0E+$0$ 94.1% N/A 104.8% N/A 91.9% N/A 37143 N/A
8.4E+01 93.9% N/A 103.7% N/A 90.9% N/A 36754 N/A
sin 5.0E+$0$ 94.9% N/A 110.2% N/A 94.6% N/A 1195 N/A
3.2E+01 79.7% N/A 96.7% N/A 72.0% N/A 22293 N/A
sqrt 8.4E+01 81.2% N/A 81.3% N/A 77.7% N/A 400 N/A
7.1E+03 62.6% N/A 62.9% N/A 56.4% N/A 363 N/A
Average of all 89.3% N/A 94.6% N/A 87.4% N/A 7531 N/A
@@ V-B2 Experiments on EPFL Benchmarks
To show the scalability of our ALS flow, we test it on the large EPFL arithmetic benchmarks. The MUSCAT method cannot handle them, so we do not compare our method with MUSCAT and we only compare our method with MECALS. Similarly, the MaxED bounds are set to $\lfloor 2^{O/10}\rfloor$ and $\lfloor 2^{O/5}\rfloor$ for each benchmark, where $O$ is the PO number of the benchmark. To accelerate our flow, we first apply the constant LACs to quickly simplify the circuit, followed by the SASIMI LACs for further simplification. Specifically, the constant LACs are first applied to the circuit POs (like truncation) until the MaxED bound is reached and then to the internal nodes of the circuit until the MaxED bound is reached. Finally, the SASIMI LACs are applied until the MaxED bound is reached. We emphasize that despite the above modification, the ALS flow is still based on the general framework presented in Fig. 2. This modification is introduced only to reduce runtime on large benchmarks, where directly considering all constant and SASIMI LACs in each iteration is computationally expensive (>> 24 hours). With the modification, we reduce the number of LACs to be checked while still following the same ALS flow in Fig. 2.
@@
Table III compares the area ratio, delay ratio, power ratio, and runtime of our flow and MECALS on the EPFL arithmetic benchmarks under the MaxED constraint. We can see that our method can handle all benchmarks in the table with an average runtime of $7531$ seconds, while MECALS cannot handle the benchmarks log2, sin, sqrt, and square (under $5.1\times 10^{7}$ MaxED bound) in 24 hours. For the benchmarks that both our method and MECALS can handle (the top part of the table), our method further reduces the area by $2.9\%$ and power by $1.9\%$ on average with a small delay overhead, while being $23.5\times$ faster than MECALS on average. Notably, for the benchmark add128, under the two MaxED bounds of $7.6\times 10^{3}$ and $5.8\times 10^{7}$, our method speeds up by $65.8\times$ and $296.7\times$, respectively, compared to MECALS and reduces area by $7.3\%$ and $6.5\%$, respectively, and power by $5.4\%$ and $4.1\%$, respectively. Only for the benchmark bar under the MaxED bound of $5.1\times 10^{7}$, our method is slightly worse than MECALS in terms of the area ratio. However, our method is far more efficient than MECALS with a competitive delay ratio. Moreover, across the cases where both methods succeed (the top part of Table III), our method finishes after an average of 3.8 iterations, and the average number of LACs applied per iteration is 11.4. In contrast, MECALS needs an average of 37.6 iterations with only one LAC applied per iteration. This again shows the efficiency of our method.
V-C Experiments Under the MaxHD Constraint
@@
TABLE IV: Comparison of our ALS flow with the MECALS method on the ISCAS and EPFL arithmetic benchmarks under the MaxHD constraint. Bold entries indicate smaller area, delay, and power ratios, or shorter runtime. N/A means that MECALS cannot obtain the final approximate circuit in 24 hours. OOM means that MECALS terminates due to out of memory.
@@
Circuit MaxHD bound Area ratio Delay ratio Power ratio Runtime/s
Ours MECALS Ours MECALS Ours MECALS Ours MECALS
c880 2 88.3% 97.3% 100.5% 105.6% 85.9% 97.6% 0.9 1.9
5 67.5% 92.4% 113.1% 106.8% 55.6% 95.3% 1.1 2.8
c1355 3 10.1% 10.4% 3.4% 3.4% 8.1% 8.3% 0.6 73
6 9.1% 10.8% 3.4% 3.4% 7.3% 8.6% 0.4 17
c1908 2 89.9% 97.3% 89.5% 101.1% 89.7% 101.0% 1.0 5.0
5 73.6% 65.5% 84.3% 86.6% 62.2% 75.3% 1.5 53
c2670 14 64.2% 84.9% 62.9% 99.6% 66.2% 83.4% 29 26
28 42.3% 41.2% 36.9% 54.4% 42.8% 44.9% 32 80
c3540 2 93.8% 95.5% 92.8% 109.1% 94.2% 97.6% 1.3 64
4 92.1% 95.6% 98.7% 102.8% 91.8% 94.6% 1.6 105
c5315 12 92.7% 96.5% 104.2% 103.5% 94.0% 96.8% 25 158
24 84.6% 90.5% 93.8% 99.6% 85.7% 91.7% 34 877
c7552 10 41.2% 95.8% 77.6% 99.8% 40.6% 96.2% 16 193
21 32.7% 84.2% 68.2% 100.1% 29.6% 83.9% 29 627
12 92.0% 92.4% 113.2% 96.3% 89.3% 90.2% 8.2 9.7
add128 25 83.7% 88.9% 101.7% 101.3% 80.4% 84.8% 35 16
12 97.0% 99.5% 101.0% 98.8% 94.5% 99.2% 27 13761
bar 25 90.8% 98.5% 99.8% 95.8% 85.0% 95.0% 35 18128
3 100.0% OOM 100.0% OOM 100.0% OOM 22045 OOM
log2 6 100.0% OOM 100.0% OOM 100.0% OOM 24133 OOM
13 86.6% 94.2% 74.7% 83.1% 82.7% 89.9% 260 38
max 26 77.0% 92.7% 64.7% 80.3% 71.7% 86.7% 172 79
12 95.9% 99.1% 98.9% 98.9% 97.1% 100.9% 357 13932
mult64 25 94.9% 98.6% 95.2% 97.9% 96.4% 100.2% 2192 22736
2 93.6% 97.9% 104.9% 101.5% 93.7% 97.9% 3002 1436
sin 5 92.8% 97.1% 100.9% 102.0% 92.7% 96.9% 4832 2062
6 81.2% N/A 82.2% N/A 78.1% N/A 560 N/A
sqrt 12 65.2% N/A 64.8% N/A 58.9% N/A 753 N/A
12 92.9% 99.0% 95.1% 95.5% 94.7% 98.4% 284 37073
square 25 91.8% 98.5% 93.4% 92.4% 92.6% 97.7% 2075 45644
Average w/o log2 & sqrt 75.8% 85.2% 83.6% 89.2% 74.0% 85.1% 517 6046
This set of experiments approximates the ISCAS85 and EPFL arithmetic benchmarks in Table I under the MaxHD constraint. We compare our ALS flow with MECALS. We modify the open-source codes of MECALS to support the MaxHD constraint and run both methods on the same platform for fair comparison. MUSCAT is not compared in this experiment as its open-source code does not support the MaxHD constraint. The MaxHD bounds are set to $\lfloor O/10\rfloor$ and $\lfloor O/5\rfloor$ for each benchmark, where $O$ is the number of POs of the benchmark. For the max benchmark, we set the parameter $K$ in the sorting strategy in Section IV-C3 to $1000$ to achieve better approximate circuits with lower hardware costs, while for the others, $K$ still keeps the default value of $100$. Similar as in Section V-B2, to accelerate our flow, the constant LACs are first applied until the MaxHD bound is reached, and then the SASIMI LACs are applied until the MaxHD bound is reached.
[…]
V-D Effectiveness of Simulation and Parameter Study
This section evaluates the effectiveness of simulation in our ALS flow. It also studies the impact of several important parameters in our method on the synthesis quality and runtime.